OpenStudy (jenniferjuice):

Solve for x: 3x − 5 = 2x + 6

ganeshie8 (ganeshie8):

To solve for x, you need to isolate x

ganeshie8 (ganeshie8):

3x − 5 = 2x + 6 subtract 2x both sides x - 5 = 6

ganeshie8 (ganeshie8):

next, add 5 both sides x = 11

OpenStudy (jenniferjuice):

im so confused

OpenStudy (jenniferjuice):

i got -1

ganeshie8 (ganeshie8):

i got 11

ganeshie8 (ganeshie8):

show ur steps :)

OpenStudy (jenniferjuice):

3x-5=2x+6 -3x -3x ------------ -5=-1x+6 -5=5x 5/-5

ganeshie8 (ganeshie8):

3x-5=2x+6 -2x -2x ------------ ?

ganeshie8 (ganeshie8):

you need to subtract 2x both sides, not 3x

OpenStudy (jenniferjuice):

ohhhhhhhhhh

ganeshie8 (ganeshie8):

:) do it and see wat u get :)

OpenStudy (jenniferjuice):

after i subract 2x from each side i got x-5=6.

ganeshie8 (ganeshie8):

Good :)

OpenStudy (jenniferjuice):

then i subtract -5 from each side and i got x=11

OpenStudy (jenniferjuice):

thankyou :)

ganeshie8 (ganeshie8):

np :)